OpenStudy (anonymous):

which is the equation with slope 5 that contains point (-2,-3) a.y-3=5(x-2) by+2=5(x-3) c.y+3=5(x+2) d.y-2=5(x-3)

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow y = 5x + b }\) Let's substitute. \(\Large \color{MidnightBlue}{\Rightarrow -3 = 5(-2) + b }\) \(\Large \color{MidnightBlue}{\Rightarrow -3 = -10 + b }\) \(\Large \color{MidnightBlue}{\Rightarrow 7 = b }\) So, the equation is: y = 5x + 7

jimthompson5910 (jim_thompson5910):

Those previous equations are in slope intercept form. The answer choices are in point-slope form (which is arguably easier to deal with since you only take one step really) y - y1 = m(x-x1) y - (-3) = 5(x - (-2) ) .... Plug in x = -2, y = -3 and m = 5 y + 3 = 5(x+2) ... Rewrite x-(-2) as x+2 and y-(-3) as y+3 So the equation in point-slope form is y + 3 = 5(x+2)

OpenStudy (anonymous):

thatnkk youu

jimthompson5910 (jim_thompson5910):

yw